OpenStudy (anonymous):

A dry cell of EMF 1.5 volt and internal resistance 0.10 ohm is connected across a resistor in series with a very low resistance ammeter.when the circuit is swithched on the ammeter reading settles to steady value of 2A.what is the (a) rate of chemical change consumption of the cell (b) rate of energy dissipation inside the cell. (c) rate of energy dessipation inside a resistor.

OpenStudy (anonymous):

@shruti you here ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

http://www.ausetute.com.au/ratelaw.html try the link for the first part i m kinda missed you right now.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@annas link is of chem. and my question is of physics.

OpenStudy (anonymous):

ok i got it E*I = rate of chemical change consumption of the cell

OpenStudy (anonymous):

ok we have I=2A V=1.5 volt R=0.10 ohm for the parts C energy dassipation inside resistor is given by P=I^2R=(2)^2(0.10)=0.4 watt

OpenStudy (anonymous):

1.5 * 2= 3 watt

OpenStudy (anonymous):

second part is done by sami

OpenStudy (anonymous):

3) I= E/(R+r)

OpenStudy (anonymous):

now for the part a rate of consumption should be equal to power output ( p=IE=1.5*2=3W

OpenStudy (anonymous):

2= 1.5/R+0.10

OpenStudy (anonymous):

R= 0.65 ohm

OpenStudy (anonymous):

@shruti the link was pretty similar to the question i know it was chemistry related i thought it would help :)

OpenStudy (anonymous):

okay, let me check it once.

OpenStudy (anonymous):

< signifies?

OpenStudy (anonymous):

as anas found R=0.65 so rate of energy dessipation inside a resistor is P=I^2R=(2)^2*0.65=2.6 Watt

OpenStudy (anonymous):

@sami-21 you're genius man...

OpenStudy (anonymous):

what does "<" signifies?

OpenStudy (anonymous):

'<' is a sign of inequality i've no idea what it signifies here

OpenStudy (anonymous):

'<'it means less than where it is mentioned here?

OpenStudy (anonymous):

P=I^2R=(2)^2*0.65=2.6 Watt

OpenStudy (anonymous):

"^" is for power

OpenStudy (anonymous):

it means \[p=i^2R\] it means 2 is in the power of I

OpenStudy (anonymous):

exponent